OpenStudy (anonymous):

Let f(x) = 4x - 5 and g(x) = 6x - 3. Find f(x) g(x). A. 10x – 8 B. 10x – 2 C. –2x – 8 D. –2x – 2

OpenStudy (anonymous):

It's A

OpenStudy (anonymous):

are you positive? can you tell me how you did it?

OpenStudy (anonymous):

I am. And I can't remember how to do it. I googled it lol

OpenStudy (anonymous):

i googled it too but all i got was the answer could be a or d.

OpenStudy (anonymous):

f(x) * g(x) = (4x - 5)(6x - 3) = 24x^2 - 12x - 30x + 15 = 24x^2 - 42x + 15

OpenStudy (anonymous):

That should explain it.

OpenStudy (anonymous):

thank you!!:)

OpenStudy (anonymous):

Np!(: